# On the Impact of Uncertainty and Calibration on Likelihood-Ratio Membership Inference Attacks

### v3

@@ I-B State of the Art
Reference [2] introduced a type of likelihood ratio attack (LiRA), in which Gaussian distributions are fitted to the TLC generated by shadow models trained to imitate models trained with and without the candidate data point. LiRA is then conducted through a parametric likelihood ratio test. Reference [16] improved LiRA by adding adversarial noise to the input in order to amplify the output gap between the two classes of shadow models. Further enhancements include setting sample-specific thresholds [17] and the use of pairwise likelihood ratio tests calibrated through a Bayesian method [18].
@@
Notably, a line of work is motivated by the opposite nature of the goal of MIAs – revealing a training data point – and of differential privacy (DP) mechanisms – obscuring the presence or absence of a training data point [30]. Based on this observation, upper bounds on the attacker’s advantage [3, 4] and positive predictive value [13] were studied for DP-protected models. These works focus on the theoretical relationship between the DP metrics and MIA success measures. Furthermore, reference [26] established universal bounds on the success rate of MIAs as a function of the target model’s generalization gap, while assuming a white-box attacker. Results that target symmetric and redundancy-invariant algorithms were also discussed in [27]. Further theoretical analyses have explored the vulnerability to MIA across population subgroups [28], as well as the impact of generalization [29].
@@ I-C Contributions and Organization
In this paper, we introduce a unified theoretical framework to analyze the performance of MIA under CV, TLC, and DS disclosure modes, assuming that the attacker follows the membership inference game studied in [3, 13, 2, 17, 18] (see Definition 1 in Sec. II-A). Rather than proposing a new attack, our objective is to provide a theoretical analysis of existing LiRA-style attacks, focusing on the impact and interplay of calibration and predictive uncertainty. The main contributions of this paper are summarized as follows.
  • As illustrated in Fig. 1, treating MIA design as a hypothesis-testing problem, we introduce a general information-theoretical framework to analyze the attacker’s advantage for CV, TLC, and DS disclosure settings. Specifically, as seen in Fig. 2, our methodology analyzes LiRA-style attacks [2] for a wide range of disclosure settings, beyond the case of TLC disclosure considered in prior work. LiRA-style attacks use shadow models to imitate the behavior of the target model for inputs inside and outside the training set.
@@ IV-D Decision Set Disclosure
The proportionality factor $\delta_{T,q}$ depends on the temperature parameter $T$ as illustrated in Fig. 4 for the case $K=2$. In particular, for $T\rightarrow 0$, which corresponds to the deterministic function in (5), the constant attains the maximum value of $1$, except for the trivial cases $q=0$ and $q=1$, which yield $\delta_{T,q}=0$ as $T\rightarrow 0$. Thus, for a deterministic disclosure function, the upper bound (40) is not sufficiently tight to quantify any decrease in the attacker’s advantage. However, for any $T>0$, and thus even with minimal randomness, the bound predicts that the attacker’s advantage at first increases as $q$ grows, since a larger threshold in (36) allows to capture more information about the confidence vector $q$. However, an excessively large $q$ ends up decreasing the performance of the attacker as the predicted set $O(p)$ in (36) gets increasingly smaller. Furthermore, as $T\rightarrow\infty$, the sigmoid function equals $\sigma_{T}(x)=1/2$ and the proportionality factor equals zero, $\delta_{T,q}=0$. In this case, the attacker does not obtain any useful information, resulting in a vanishing advantage.
@@ V-A On the Dirichlet Assumption
First, to validate the use of a Dirichlet distribution to model the confidence probability, we evaluate the true distribution of the confidence vector produced by a neural network-based classifier. Following the setting in [1], we use the CIFAR-10 data set, a standard benchmark for image classification involving images classified by $K=10$ labels. To construct the training sets $\mathcal{D}^{\mathrm{out}}$ and $\mathcal{D}^{\mathrm{in}}$, we draw $N^{\mathrm{tr}}=4000$ samples at random from the overall data set. The target classification model adopts a standard convolutional neural network (CNN) with two convolution and max pooling layers, followed by a $128$-unit fully connected layer and a SoftMax output layer. The Tanh is used as the activation function, and the cross-entropy serves as the loss function. We adopt the stochastic gradient descent (SGD) optimizer with a learning rate of $0.001$, momentum of $0.9$, and weight decay of $1\times 10^{-7}$, over $100$ epochs for model training.
@@ VI Conclusions
Given that the derived bounds assume a worst-case adversary, the analysis in this paper is also applicable to recent LiRA-style atacks that inherently limit attack performance. These include settings with data shifts unknown to the attacker, as well as computational constraints [16, 17, 18]. Notably, reference [18] uses a Bayesian method to provide a lightweight version of LiRA-style attack, requiring only one or two shadow models.
@@
  • Calibration error: Regularization-based defenses, such as L2 regularization [1, 20], dropout [49], pruning [50], and model stacking [11], reduce overfitting, which helps lower the model’s confidence in its predictions. This makes the model more robust to MIAs by diminishing the risk of overconfidence, which attackers can exploit. Furthermore, to directly reduce calibration error, methods such as temperature scaling, Platt scaling, and isotonic regression [34] can be used to reduce the model’s overconfidence.
@@
To analyze the underlying causes of the above empirical observations within our theoretical framework, we conduct experiments on the MNIST data set, treating class 00 as a distinct subgroup and classes $1$ to $9$ as the majority group. The classification model is a convolutional neural network with two convolutional layers ($32$ and $64$ filters) followed by two fully connected layers ($128$ and $10$ units). We train the model using the Adam optimizer with a learning rate of $0.001$ for $20$ epochs, minimizing cross-entropy loss. To simulate the aleatoric uncertainty, we set the ground-truth probability for class 00 to $p^{*}_{0}=0.8$ and introduce label noise as in the previous example. The relative calibration error is then evaluated using equation (24) from our paper.

### v4

@@ I-B State of the Art
Reference [2] introduced a type of likelihood ratio attack (LiRA), in which Gaussian distributions are fitted to the TLC generated by shadow models trained to imitate models trained with and without the candidate data point. LiRA is then conducted through a parametric likelihood ratio test. Reference [16] improved LiRA by adding adversarial noise to the input in order to amplify the output gap between the two classes of shadow models. Further enhancements include setting sample-specific thresholds [17] and the use of pairwise likelihood ratio tests calibrated through a Bayesian method [18]. We refer to this family of methods as LiRA-style attacks, which includes the original LiRA [2] and its subsequent variants [16, 17, 18], all of which follow the likelihood-ratio-based hypothesis testing paradigm using shadow models to estimate membership likelihoods.
@@
Notably, a line of work is motivated by the opposite nature of the goal of MIAs – revealing a training data point – and of differential privacy (DP) mechanisms – obscuring the presence or absence of a training data point [30]. Based on this observation, upper bounds on the attacker’s advantage [3, 4] and positive predictive value [13] were studied for DP-protected models. These works focus on the theoretical relationship between the DP metrics and MIA success measures. Furthermore, reference [26] established universal bounds on the success rate of MIAs as a function of the target model’s generalization gap, while assuming a Bayesian attacker. Results that target symmetric and redundancy-invariant algorithms were also discussed in [27]. Further theoretical analyses have explored the vulnerability to MIA across population subgroups [28], as well as the impact of generalization [29].
Compared to these prior theoretical analyses, our work differs in four key aspects: 1) Most existing studies focus on TLC disclosure, typically in the form of losses [3, 4, 13, 26, 28, 29]. In contrast, we systematically analyze three widely-used disclosure modes, i.e., CV, TLC, and DS, which have not been jointly studied in theoretical work. 2) Prior analyses adopt fundamentally different frameworks, such as DP-specific [3, 4, 13] and Bayesian [26] settings. Our framework is specifically designed for LiRA-style attacks [2, 16, 17, 18], a class of practical and efficient likelihood-ratio-based MIAs that lack theoretical attention. 3) While some previous works often aim to establish tight bounds [3, 4, 13], our goal is to derive trend-aware bounds that reveal how key factors influence MIA performance, prioritizing interpretability over tightness. 4) Finally, prior studies focus on factors such as model architecture [4], data size [27] and training epochs [29]. We instead focus on deeper underlying factors, i.e., aleatoric uncertainty, epistemic uncertainty, and calibration error, which, as shown in Appendix -G, explain and unify many lower-level effects observed in prior work [1, 11, 24, 28].
@@ I-C Contributions and Organization
In this paper, we introduce a unified theoretical framework to analyze the performance of MIA under CV, TLC, and DS disclosure modes, assuming that the attacker follows the membership inference game studied in [3, 13, 2, 17, 18] (see Definition 1 in Sec. II-A). Rather than proposing a new attack, our objective is to provide a theoretical analysis of existing LiRA-style attacks [2, 16, 17, 18], focusing on the impact and interplay of calibration and predictive uncertainty. The main contributions of this paper are summarized as follows.
  • As illustrated in Fig. 1, treating MIA design as a hypothesis-testing problem, we introduce a general information-theoretical framework to analyze the attacker’s advantage for CV, TLC, and DS disclosure settings. Specifically, as seen in Fig. 2, our methodology analyzes LiRA-style attacks [2, 16, 17, 18] for a wide range of disclosure settings, beyond the case of TLC disclosure considered in prior work. LiRA-style attacks use shadow models to imitate the behavior of the target model for inputs inside and outside the training set.
@@ IV-D Decision Set Disclosure
The proportionality factor $\delta_{T,q}$ depends on the temperature parameter $T$ as illustrated in Fig. 4 for the case $K=2$. In particular, for $T\rightarrow 0$, which corresponds to the deterministic function in (5), the constant attains the maximum value of $1$, except for the trivial cases $q=0$ and $q=1$, which yield $\delta_{T,q}=0$ as $T\rightarrow 0$. Thus, for a deterministic disclosure function, the upper bound (40) is not sufficiently tight to quantify any decrease in the attacker’s advantage. However, for any $T>0$, and thus even with minimal randomness, the bound predicts that the attacker’s advantage at first increases as $q$ grows, since a larger threshold in (36) allows to capture more information about the confidence vector $p$. However, an excessively large $q$ ends up decreasing the performance of the attacker as the predicted set $O(p)$ in (36) gets increasingly smaller. Furthermore, as $T\rightarrow\infty$, the sigmoid function equals $\sigma_{T}(x)=1/2$ and the proportionality factor equals zero, $\delta_{T,q}=0$. In this case, the attacker does not obtain any useful information, resulting in a vanishing advantage.
@@ V-A On the Dirichlet Assumption
First, to validate the use of a Dirichlet distribution to model the confidence probability, we evaluate the true distribution of the confidence vector produced by a neural network-based classifier. Following the setting in [1, 2, 16, 17, 18], we use the CIFAR-10 data set, a standard benchmark for image classification involving images classified by $K=10$ labels. To construct the training sets $\mathcal{D}^{\mathrm{out}}$ and $\mathcal{D}^{\mathrm{in}}$, we draw $N^{\mathrm{tr}}=4000$ samples at random from the overall data set. The target classification model adopts a standard convolutional neural network (CNN) with two convolution and max pooling layers, followed by a $128$-unit fully connected layer and a SoftMax output layer. The Tanh is used as the activation function, and the cross-entropy serves as the loss function. We adopt the stochastic gradient descent (SGD) optimizer with a learning rate of $0.001$, momentum of $0.9$, and weight decay of $1\times 10^{-7}$, over $100$ epochs for model training.
@@ VI Conclusions
Given that the derived bounds assume a worst-case adversary, the analysis in this paper is also applicable to recent LiRA-style attacks that inherently limit attack performance. These include settings with data shifts unknown to the attacker, as well as computational constraints [16, 17, 18]. Notably, reference [18] uses a Bayesian method to provide a lightweight version of LiRA-style attack, requiring only one or two shadow models.
Note that empirical evaluation under CV and DS disclosures becomes infeasible when the number of classes is large, due to the challenge of high-dimensional distribution estimation. This is an inherent limitation of the LiRA-style attack paradigm. This is why existing implementations [2, 16, 17, 18] focus exclusively on TLC disclosures. Nonetheless, our theoretical results in Propositions 13 remain valid for arbitrary class sizes and continue to offer meaningful insights across all disclosure types.
@@
  • Calibration error: Regularization-based defenses, such as L2 regularization [1, 20], dropout [49], pruning [50], and model stacking [11], reduce overfitting, which helps lower the model’s confidence in its predictions. This makes the model more robust to MIAs by diminishing the risk of overconfidence, which attackers can exploit. Furthermore, to directly reduce calibration error, methods such as temperature scaling [51], Platt scaling, and isotonic regression [34] can be used to reduce the model’s overconfidence.
@@
To analyze the underlying causes of the above empirical observations within our theoretical framework, we conduct experiments on the MNIST data set, treating class 00 as a distinct subgroup and classes $1$ to $9$ as the majority group. The classification model is a CNN with two convolutional layers ($32$ and $64$ filters) followed by two fully connected layers ($128$ and $10$ units). We train the model using the Adam optimizer with a learning rate of $0.001$ for $20$ epochs, minimizing cross-entropy loss. To simulate the aleatoric uncertainty, we set the ground-truth probability for class 00 to $p^{*}_{0}=0.8$ and introduce label noise as in the previous example. The relative calibration error is then evaluated using equation (24) from our paper.